OpenStudy (anonymous):

what is the area of enclosed region by y=1/x, y=1/x^2, x=2?

OpenStudy (mr.math):

Can you draw the region? This is the point you should start at.

OpenStudy (dumbcow):

\[\int\limits_{1}^{2}\frac{1}{x} - \frac{1}{x^{2}}dx\]